OpenStudy (anonymous):

under the QUADRATIC EQUATION(Roots) Find the number, when added to its square , the sum is 42.

OpenStudy (anonymous):

x + x^2 = 42...

OpenStudy (anonymous):

x = 6 or x = -7

OpenStudy (anonymous):

\[x+x^2=42\]\[x^2+x-42=0\]\[(x+7)(x-6)=0\]\[x=-7, 6\] If you try plugging both numbers in they work.